### v1

@@ 6 Conclusions
In this paper, we developed a relative-type inexact proximal augmented Lagrangian method (ripALM) for solving a class of linearly constrained convex optimization problems. The proposed ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. By employing a relative-type error criterion, it simplifies implementation and parameter tuning, compared to the absolute-type inexact pALM. We conducted a thorough convergence analysis and demonstrated the competitive efficiency of our ripALM through numerical experiments on solving quadratically regularized optimal transport problems. Since our algorithm is applicable to a more general linearly constrained convex optimization problem, future work may explore potential applications of this framework to other practical problems, including general regularized optimal transport problems studied in [51].
@@ Appendix A Missing Proofs in Section 3
Since $\{\nu_{k}\}$ is a non-negative summable sequence, it then follows from [32, Lemma 2 in Section 2.2] that the sequence $\left\{\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ is convergent. This together with $\tau_{k}\geq\tau_{\min}>0$ implies that all sequences $\{\bm{x}^{k}\}$, $\{\bm{w}^{k}\}$, $\{\bm{y}^{k}\}$ are bounded.
@@
Since $\{\sigma_{k}\}$ is bounded away from 00, we then obtain that $\lim\limits_{k\to\infty}\langle\bm{w}^{k}-\bm{y}^{k+1},\,\Delta^{k+1}\rangle=0$ and $\lim\limits_{k\to\infty}\Delta^{k+1}=\bm{0}$. Finally, recall again that $\tau_{k+1}\leq(1+\nu_{k})\tau_{k}$ with $\nu_{k}\geq 0$ and $\sum\nu_{i}<\infty$ for all $k\geq 0$. Thus, $\tau_{k}$ must be bounded from above and hence $\tau_{k}\sigma_{k}^{-1}$ is also bounded from above. Consequently, we can obtain that $\lim\limits_{k\to\infty}\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})=\bm{0}$ and hence $\lim\limits_{k\to\infty}\bm{\theta}^{k+1}=\bm{0}$.
@@
Note that $F$ is closed upper semicontinuous concave (see, for example, [37, Theorem 7]), and thus $\mathrm{dom}\,F$ is closed. This together with $(\bm{0},\,\bm{x}^{\infty})=\lim\limits_{i\to\infty}\big{(}\bm{\theta}^{k_{i}+1},\,\bm{x}^{k_{i}+1}\big{)}$ induces that $(\bm{0},\,\bm{x}^{\infty})\in\mathrm{dom}\,F$. Then, we see that
$\;\;\;\;\;F(0,\,\bm{x}^{*})$
@@
We next study the limit of $\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\}$. Since $-F$ and $G$ are convex conjugate and $\big{(}\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big{)}\in\partial G(\bm{y}^{k+1},\,\bm{\xi}^{k+1})$, we can get the following equality by using the Fenchel equality (see, for example, [36, Theorem 23.5]):
@@
Statement (iv). We first prove that any accumulation point of $\{\bm{y}^{k}\}$ is an optimal solution of problem (2.1). Since $\{\bm{y}^{k}\}$ is bounded by statement (i), the sequence $\{\bm{y}^{k}\}$ has at least one accumulation point. Suppose that $\bm{y}^{\infty}$ is an accumulation point and $\{\bm{y}^{k_{j}}\}$ is a convergent subsequence such that $\lim\limits_{j\to\infty}\bm{y}^{k_{j}}=\bm{y}^{\infty}$. Since $\lim\limits_{k\to\infty}\|\bm{y}^{k+1}-\bm{y}^{k}\|=0$, we also have that $\lim\limits_{j\to\infty}\bm{y}^{k_{j}+1}=\bm{y}^{\infty}$. Then, using the fact that $G$ is lower semicontinuous and convex, and $\lim\limits_{k\to\infty}\bm{\xi}^{k+1}=0$, we obtain that
@@
where $\overline{\tau}_{k}$ is defined as $\overline{\tau}_{k}:=\max\left\{1,\tau_{k}\right\}$ and the last inequality follows from (A.17). Now, combing (A.17), (A.18) and (A.19), with some manipulations, we can obtain that
@@
$\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)>\widetilde{c}:=\frac{c-1}{c}\cdot\frac{\sqrt{\tau_{\min}}-2\sqrt{\rho}}{\sqrt{\tau_{\min}}}>0,\quad\forall\,k\geq k_{0},$ (A.21)
@@
$\gamma_{k}\,:=$ $\,\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{k}}\right)^{2}\overline{\tau}_{k}}$ (A.23)
@@
$\geq$ $\,\gamma_{\min}:=\frac{\widetilde{c}\,\sigma_{\min}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{\max}}\right)^{2}\overline{\tau}_{\max}}>0,\quad\forall\,k\geq k_{0}.$
@@
where $\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}$. Since $\nu_{k}\to 0$ and $\gamma_{k}\geq\gamma_{\min}>0$ for all $k>k_{0}$, one can verify that $\limsup\limits_{k\to\infty}\,\{\mu_{k}\}<1$. Thus, we obtain the desired results in statement (i).
@@
$\widetilde{c}\,\tau_{\min}\|\bm{y}^{k+1}-{\bm{y}}^{k}\|^{2}\leq\widetilde{c}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-{\bm{y}}^{k})\\ \bm{x}^{k+1}-{\bm{x}}^{k}\end{matrix}\right\|^{2}\leq\mathrm{dist}_{\Lambda^{k}}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right),$
for any $k\geq k_{0}$. Using this inequality and the fact that the sequence $\left\{\mathrm{dist}_{\Lambda^{k}}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\right\}$ is asymptotically Q-(super)linear convergent, we can conclude that there exist a positive integer $k_{1}$, $10<\beta<$ 1 and $C>0$ such that
@@
Appendix B A dual ADMM for QROTs
@@
::=:= $\;f_{\texttt{q}}^{*}(W)-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\left\langle X,\,\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-W\right\rangle+\cfrac{\sigma}{2}\left\|\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-W\right\|_{F}^{2}.$
@@
[…]
@@
$W^{k+1}~{}:=~{}\texttt{prox}_{\sigma^{-1}f_{\texttt{q}}^{*}}\big{(}Q^{k}\big{)}=~{}\left\{\begin{aligned} &C-\Pi_{\mathbb{R}^{m\times n}_{+}}\big{(}C-Q^{k}\big{)},&&\lambda=0,\\[3.0pt] &Q^{k}-(1+\lambda\sigma)^{-1}\Pi_{\mathbb{R}^{m\times n}_{+}}(Q^{k}-C),&&\lambda>0,\end{aligned}\right.$
@@
Appendix C An iBPGM for QROTs
In this section, we briefly discuss how to employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1). We refer readers to [52, Section 5] for more details. Specifically, the iBPGM with the entropy kernel function for solving (4.1) can be given as follows: let $X^{0}>0$ and $\phi(X):=\sum_{ij}x_{ij}(\log x_{ij}-1)$, at the $k$-th iteration, compute
$X^{k+1}~{}\approx{}$ $\min\limits_{X}\Big{\{}\langle C+\lambda X^{k},\,X\rangle+\mu_{k}\mathcal{D}_{\phi}(X,\,X^{k})\;\mid\;X\bm{1}_{n}=\bm{\alpha},~{}~{}X^{\top}\bm{1}_{m}=\bm{\beta}\Big{\}},$ (C.1)
@@ Appendix C An iBPGM for QROTs
where $M^{k}:=C+\lambda X^{k}-\mu_{k}\log X^{k}$. Note that problem (C.2) has the same form as the entropic regularized optimal transport problem and hence can be readily solved by the popular Sinkhorn’s algorithm; see [31, Section 4.2] for more details. Specifically, let $\Xi^{k}:=e^{-M^{k}/\mu_{k}}$. Then, given an initial positive vector $\bm{v}^{k,0}$, the iterative scheme is given by

### v2

@@ 6 Conclusions
In this paper, we developed a relative-type inexact proximal augmented Lagrangian method (ripALM) for solving a class of linearly constrained convex optimization problems. The proposed ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. By employing a suitable relative-type error criterion, it simplifies implementation and parameter tuning, compared to the absolute-type inexact pALM. We conducted a thorough convergence analysis and demonstrated the competitive efficiency of our ripALM through numerical experiments on solving quadratically regularized optimal transport problems and basis pursuit denoising problems. Since our model covers a wide range of applications, future work may explore the potential of applying ripALM to other practical problems, including general regularized optimal transport problems as studied in [58]. Additionally, it would be interesting to extend ripALM to a broader class of convex optimization problems, thereby further enhancing its applicability.
Acknowledgements
We thank the associate editor and two anonymous reviewers for their valuable comments and suggestions, which have helped to improve the quality of this paper. The research of Lei Yang is supported in part by the National Key Research and Development Program of China under grant 2023YFB3001704 and the National Natural Science Foundation of China under grant 12301411. The research of Kim-Chuan Toh is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 grant (MOE-T2EP20224-0017).
@@ Appendix A Missing Proofs in Section 3
Since $\{\nu_{k}\}$ is a non-negative summable sequence, it then follows from [35, Lemma 2 in Section 2.2 on Page 44] that the sequence $\left\{\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ is convergent (but do not necessarily converge to 0). This together with $\tau_{k}\geq\tau_{\min}>0$ implies that all sequences $\{\bm{x}^{k}\}$, $\{\bm{w}^{k}\}$, $\{\bm{y}^{k}\}$ are bounded.
@@
Since $\{\sigma_{k}\}$ is bounded away from 0, we then obtain that $\lim\limits_{k\to\infty}\langle\bm{w}^{k}-\bm{y}^{k+1},\,\Delta^{k+1}\rangle=0$ and $\lim\limits_{k\to\infty}\Delta^{k+1}=\bm{0}$. Finally, recall again that $\tau_{k+1}\leq(1+\nu_{k})\tau_{k}$ with $\nu_{k}\geq 0$ and $\sum\nu_{i}<\infty$ for all $k\geq 0$. Thus, $\tau_{k}$ must be bounded from above and hence $\tau_{k}\sigma_{k}^{-1}$ is also bounded from above. Consequently, we can obtain that $\lim\limits_{k\to\infty}\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})=\bm{0}$ and hence $\lim\limits_{k\to\infty}\bm{\theta}^{k+1}=\bm{0}$.
@@
Note that $F$ is closed upper semicontinuous concave (see, for example, [40, Theorem 7]), and thus $\mathrm{dom}\,F$ is closed. This together with $(\bm{0},\,\bm{x}^{\infty})=\lim\limits_{i\to\infty}\big(\bm{\theta}^{k_{i}+1},\,\bm{x}^{k_{i}+1}\big)$ induces that $(\bm{0},\,\bm{x}^{\infty})\in\mathrm{dom}\,F$. Then, we see that
$\;\;\;\;\;F(\bm{0},\,\bm{x}^{*})$
@@
We next study the limit of $\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\}$. Since F-F and $G$ are convex conjugate and $\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)\in\partial G(\bm{y}^{k+1},\,\bm{\xi}^{k+1})$, we can get the following equality by using the Fenchel equality (see, for example, [39, Theorem 23.5]):
@@
Statement (iv). We first prove that any accumulation point of $\{\bm{y}^{k}\}$ is an optimal solution of problem (2.1). Since $\{\bm{y}^{k}\}$ is bounded by statement (i), the sequence $\{\bm{y}^{k}\}$ has at least one accumulation point. Suppose that $\bm{y}^{\infty}$ is an accumulation point and $\{\bm{y}^{k_{j}}\}$ is a convergent subsequence such that $\lim\limits_{j\to\infty}\bm{y}^{k_{j}}=\bm{y}^{\infty}$. Since $\lim\limits_{k\to\infty}\|\bm{y}^{k+1}-\bm{y}^{k}\|=0$, we also have that $\lim\limits_{j\to\infty}\bm{y}^{k_{j}+1}=\bm{y}^{\infty}$. Then, using the fact that $G$ is lower semicontinuous and convex, and $\lim\limits_{k\to\infty}\bm{\xi}^{k+1}=\bm{0}$, we obtain that
@@
where $\overline{\tau}_{k}$ is defined as $\overline{\tau}_{k}:=\max\left\{1,\tau_{k}\right\}$ and the last inequality follows from (A.17). Now, combining (A.17), (A.18) and (A.19), with some manipulations, we can obtain that
@@
$\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)>\widetilde{c}:=\Big(\frac{c-1}{c}\Big)\cdot\frac{\sqrt{\tau_{\min}}-2\sqrt{\rho}}{\sqrt{\tau_{\min}}}>0,\quad\forall\,k\geq k_{0},$ (A.21)
@@
$\gamma_{k}\,$= $\,\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{k}}\right)^{2}\overline{\tau}_{k}}$ (A.23)
@@
$\geq$ $\,\gamma_{\min}=\frac{\widetilde{c}\,\sigma_{\min}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{\max}}\right)^{2}\overline{\tau}_{\max}}>0,\quad\forall\,k\geq k_{0}.$
@@
where $\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}$. Since $\nu_{k}\to 0$ and $\gamma_{k}\geq\gamma_{\min}>0$ for all $k\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\geq k_{0}$, one can verify that $\limsup\limits_{k\to\infty}\,\{\mu_{k}\}<1$. Thus, we obtain the desired results in statement (i).
@@
$\widetilde{c}\,\tau_{\min}\|\bm{y}^{k+1}-{\bm{y}}^{k}\|^{2}\leq\widetilde{c}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-{\bm{y}}^{k})\\ \bm{x}^{k+1}-{\bm{x}}^{k}\end{matrix}\right\|^{2}\leq\mathrm{dist}_{\Lambda^{k}}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}2}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right),$
for any $k\geq k_{0}$. Using this inequality and the fact that the sequence $\left\{\mathrm{dist}_{\Lambda^{k}}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\right\}$ is asymptotically Q-(super)linear convergent, we can conclude that there exist a positive integer $k_{1}$, $0<\beta<1$ and $C>0$ such that
@@
Appendix B A dual ADMM for QROT
@@
= $\;f_{\texttt{q}}^{*}(W)-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\big\langle X,\,\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-W\big\rangle+\frac{\sigma}{2}\big\|\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-W\big\|_{F}^{2}.$
@@
$\bm{u}^{*}(t)$ $=$ $\frac{\sigma^{-1}\bm{\alpha}+S^{k}\bm{1}_{n}}{n}+t\,\bm{1}_{m},\quad\forall\,t\in\mathbb{R},$
@@
$W^{k+1}~:=~\mathtt{prox}_{\sigma^{-1}f_{\texttt{q}}^{*}}\big(Q^{k}\big)=~\left\{\begin{aligned} &C-\Pi_{\mathbb{R}^{m\times n}_{+}}\big(C-Q^{k}\big),&&\lambda=0,\\[3.0pt] &Q^{k}-(1+\lambda\sigma)^{-1}\Pi_{\mathbb{R}^{m\times n}_{+}}(Q^{k}-C),&&\lambda>0,\end{aligned}\right$.
@@
Appendix C An iBPGM for QROT
In this section, we briefly discuss how to employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1). We refer readers to [59, Section 5] for more details. Specifically, the iBPGM with the entropy kernel function for solving (4.1) can be given as follows: let $X^{0}>0$ and $\phi(X):=\sum_{ij}x_{ij}(\log x_{ij}-1)$, at the $k$-th iteration, compute
$X^{k+1}$$\approx$ $\min\limits_{X}\Big\{\langle C+\lambda X^{k},\,X\rangle+\mu_{k}\mathcal{D}_{\phi}(X,\,X^{k})\;\mid\;X\bm{1}_{n}=\bm{\alpha},~~X^{\top}\bm{1}_{m}=\bm{\beta}\Big\},$ (C.1)
@@ Appendix C An iBPGM for QROT
where $M^{k}:=C+\lambda X^{k}-\mu_{k}\log X^{k}$. Note that problem (C.2) has the same form as the entropic regularized optimal transport problem and hence can be readily solved by the popular Sinkhorn’s algorithm; see [34, Section 4.2] for more details. Specifically, let $\Xi^{k}:=e^{-M^{k}/\mu_{k}}$. Then, given an initial positive vector $\bm{v}^{k,0}$, the iterative scheme is given by